# Line Cover and Related Problems

### v1

@@ Abstract
Our main contributions reveal interesting differences in the parameterized complexity of these problems. While Line Cover is fixed-parameter tractable parameterized by the number $k$ of lines in the solution, we show that Line Clustering is $\operatorname{W}[1]$-hard when parameterized by $k$ and rule out algorithms of running time $n^{o(k)}$ under the Exponential Time Hypothesis. Hyperplane Cover is known to be NP-hard even when $d=2$ and by the work of Langerman and Morin [Discrete & Computational Geometry, 2005], it is $\operatorname{FPT}$ parameterized by $k$ and $d$. We complement this result by establishing that Hyperplane Cover is $\operatorname{W}[2]$-hard when parameterized by only $k$.
We complement our hardness results by presenting an algorithm for Projective Clustering. We show that this problem is solvable in $n^{\mathcal{O}(dk(r+1))}$ time. Not only does this yield an upper bound for Line Clustering that asymptotically matches our lower bound, but it also significantly extends the seminal work on $k$-Means Clustering (the special case $r=0$) by Inaba, Katoh, and Imai [SoCG 1994].
@@ 1 Introduction
The Line Cover problem asks for the smallest number $k$ of lines that cover a given set of $n$ points in the Euclidean plane $\mathbb{R}^{2}$. This problem is NP-hard and is well-studied in the context of parameterized complexity, particularly because it admits a simple kernelization, often highlighted in introductory lectures on kernelization [cygan2015parameterized, fomin2019kernelization, Niedermeierbook06].
@@
Line Clustering is an optimization variant of Line Cover, in which the goal is to find $k$ lines that best fit the points in the $\ell_{2}$ norm, rather than merely covering them. In other words, we search for $k$ lines such that the sum of the squares distances from each point to the closest point on one of the lines is minimized.
@@
Figure 1 provides an example of Line Clustering with $k=3$. In Operations Research, this problem is known as Line Facility Location  [Megiddo1982]. The problem has been studied for convex and piecewise linear functions [BOBROWSKI2012108] as well as in the weighted setting [DBLP:journals/jco/CheungD11].
Figure 1: An instance of Line Clustering with $k=3$ and a solution (the three colored lines). The points are colored based on the color of the nearest line in the selected solution. The dashed line show the region where points have equal distance from two solution lines.
@@
Langerman and Morin [langerman2005covering] initiated the study of the parameterized complexity of Hyperplane Cover (as part of a more general algorithmic framework). They showed that the problem is FPT when parameterized by $k+d$, providing several algorithms running in $k^{\mathcal{O}(dk)}n^{\mathcal{O}(1)}$ time.
@@
Projective Clustering naturally finds many applications in fields such as unsupervised learning [DBLP:reference/ml/Procopiuc10], data mining [AggarwalPWYP99, parsons2004subspace], artificial intelligence [10460289], computational biology [DBLP:reference/ml/Procopiuc10], database management [DBLP:conf/vldb/ChakrabartiM00], computer vision [DBLP:conf/sigmod/ProcopiucJAM02], image segmentation [yang2008unsupervised], motion segmentation [vidal2008multiframe], face clustering [ho2003clustering], image processing [hong2006multi], systems theory [vidal2003algebraic], and others [elhamifar2013sparse, vidal2011subspace]. It also generalizes a number of fundamental and well-studied problems such as $k$-Means Clustering [Lloyd82], Principal Component Analysis (PCA[Eckart1936TheAO], and Hyperplane Approximation [korneenko1993hyperplane].
@@
We consider 1 to be particularly interesting for several reasons. First, it underscores a clear contrast between Line Cover—which is $\operatorname{FPT}$ when parameterized by kk—and Line Clustering. Second, Line Clustering is closely related to $k$-means clustering, where instead of fitting $k$ lines, one seeks $k$ points (i.e., zero-dimensional subspaces) that best fit the data in the $L_{2}$ norm. Notably, the parameterized complexity of $k$-means clustering in $\mathds{R}^{2}$ (with parameter $k$) remains a longstanding open question [cohen-addad2018bane].
Our second lower bound concerns Hyperplane Cover. Since in $\mathds{R}^{2}$ this problem is equivalent to Line Cover, which is $\operatorname{NP}$-complete, we know that the problem is $\operatorname{Para-NP}$-complete when parameterized by $d$. On the other hand, by the work of Langerman and Morin [langerman2005covering], Hyperplane Cover is $\operatorname{FPT}$ when parameterized by $k$ and $d$. 2 refines this complexity picture by proving that Hyperplane Cover is $\operatorname{W}[2]$-hard when parameterized by $k$ alone (and $d$ is large compared to $k$).
@@
Hyperplane Cover is $\operatorname{W}[2]$-hard parameterized by $k$. Assuming the ETH, it cannot be solved in $n^{o(k)}$ time.
@@
Since $k$-Means Clustering is the (very) special case of Projective Clustering with $r=0$, 3 significantly extends the work of Inaba et al. [inaba1994applications], who gave an algorithm for $k$-Means Clustering running in $n^{\mathcal{O}(kd)}$ time. It is also worth noting that the tools we use to prove 3 are quite different from those employed by Inaba et al., who relied on weighted Voronoi diagrams. In contrast, the proof of 3 leverages a fundamental result from computational algebraic geometry, which computes a set of sample points that meets every realizable sign condition of a family of polynomials restricted to an algebraic set.
@@
Line Cover is NP-hard and APX-hard [kumar2000hardness, broden2001guarding, Megiddo1982]. A number of results in parameterized complexity and kernelization in the literature is devoted to Line Cover and Hyperplane Cover. Langerman and Morin [langerman2005covering] proposed an $k^{dk+d}n^{\mathcal{O}(1)}$-time algorithm for Hyperplane Cover. Later, the running time was improved by Wang et al. [wang2010parameterized] to $\tfrac{k^{(d-1)k}}{1.3^{k}}n^{\mathcal{O}(1)}$. Afshani et al. [AfshaniBDN16] gave an algorithm of running time $O(2^{n})$. They also improved on the $\operatorname{FPT}$ algorithm for Hyperplane Cover in $\mathbb{R}^{3}$. Their improved algorithm runs in $\left(\tfrac{Ck^{2}}{\log k^{1/5}}\right)^{k}n^{\mathcal{O}(1)}$ time. They also presented a $\left(\tfrac{Ck}{\log k}\right)^{k}n^{\mathcal{O}(1)}$-time algorithm for Line Cover. Kratsch et al. show that the well-known $\mathcal{O}(k^{2})$ kernel for Line Cover is tight under standard complexity assumptions [KratschPR14]. A similar result exists for Hyperplane Cover [BoissonnatDGK18].
Due to its importance in theory as well as applications, the study of Projective Clustering enjoys a long and rich history. Various approaches have been used to solve Projective Clustering: kernel sets [DBLP:journals/jacm/AgarwalHV04, DBLP:journals/siamcomp/Har-PeledW04], coresets [DBLP:journals/siamcomp/Har-PeledW04, DBLP:conf/fsttcs/VaradarajanX12], total sensitivity [DBLP:conf/stoc/FeldmanL11, DBLP:conf/soda/LangbergS10, DBLP:conf/fsttcs/VaradarajanX12], and dimension reduction via sampling [DBLP:journals/toc/DeshpandeRVW06, DBLP:journals/dcg/ShyamalkumarV12]. There is also a significant body of work on coresets constructions [DBLP:conf/stoc/DeshpandeV07, har2004coresets, DBLP:conf/stoc/HuangV20] and dimension reductions [CharikaW25] for Projective Clustering in computational geometry.
@@ 2 Preliminaries
It is widely believed that $\operatorname{FPT}\neq\operatorname{W}[1]$, and hence, if a problem is hard for a class $\operatorname{W}[i]$ with $i\geq 1$, then it is considered to be fixed-parameter intractable. For a detailed introduction to parameterized complexity, we refer readers to the textbooks by Cygan et al. and by Downey and Fellows [cygan2015parameterized, DowneyF99]. We also provide conditional lower bounds by making use of the following complexity hypothesis formulated by Impagliazzo, Paturi, and Zane [ImpagliazzoPZ01].
@@ 3 W[1]-hardness of projective clustering in the Euclidean plane
We give a reduction from Regular Multicolored Independent Set. Therein, one is given a graph $G=(V,E)$ and a partition of the $n$ vertices into $V_{1},V_{2},\ldots,V_{\ell}$ (often referred to as colors of vertices) such that $|V_{i}|=\nu=\nicefrac{{n}}{{\ell}}$ for each $i\in[\ell]$ and each vertex $v$ has exactly $q$ neighbors (none of which have the same color as $v$). The question is whether there exists an independent set containing exactly one vertex of each color. It is known that Regular Multicolored Independent Set is $\operatorname{W}[1]$-hard parameterized by $\ell$ [cygan2015parameterized, Proposition 13.5.] and cannot be solved in $n^{o(\ell)}$ time unless the $\operatorname{ETH}$ fails [Chen20061346]. We will assume without loss of generality that $\nu$ is divisible by $4$, $\nu>\ell^{3}$, $\ell>10$, and $V_{i}=\{w_{1}^{i},w_{2}^{i},\ldots,w_{\nu}^{i}\}$.
@@
In the remaining part of the proof, we provide the gap-reduction for a multiset of input points. Before we present the reduction in detail, we first give a high-level overview. Our aim is to build a grid-like structure where any optimal solution contains $\ell$ vertical and $\ell$ horizontal lines. Each horizontal line and each vertical line correspond to picking a vertex in a solution. The aim is then to construct a set of points such that (i) for each horizontal line in the solution there is a vertical line in the solution that encodes the same vertex (to ensure that a solution selects precisely $\ell$ vertices) and (ii) for each pair of horizontal and vertical lines in a solution, the corresponding vertices are not adjacent in the input graph. We do this by placing points on the intersection of two such lines whenever the two respective vertices share an edge or are two distinct vertices of the same color. Otherwise, we place a point on the horizontal line close to the intersection. Imagine for now that these points have distinct distances from a “solution line” such that no vertical line contains more than one such point. See fig. 2 for an example.
@@
With this intuition in mind, there are two main obstacles to overcome. First, we want to ensure that the solution lines are actually (almost) horizontal and vertical and second, covering a maximum number of points is different from minimizing the sum of squares of distances from each point to the closest line. We start with the second obstacle. Assume that there are two lines (one vertical and one horizontal) that is part of any optimal solution that is far enough away from the previously described set of points so that they are never the closest line to any of these points. Let $W$ be some sufficiently large (but also not too large) number. We place all vertical lines corresponding to vertices within the same set $V_{i}$ closer together than lines corresponding to vertices in different sets (as already indicated in fig. 2). We do the same for all horizontal lines. By making the distance $d_{s}$ between lines corresponding to vertices in different sets large enough, we can ensure that for any set $V_{i}$ any optimal solution contains one horizontal line and one vertical line that corresponds to a vertex in $V_{i}$ (assuming that all solution lines are horizontal or vertical). Next, we compute for each horizontal line the sum of squares of distances of all points on horizontal lines corresponding to vertices within the same set $V_{i}$ except for those points that lie on the vertical line corresponding to the same vertex. For a line $h$, let that number be $\phi(h)$. Then, we place $W-\phi(h)$ points on line $h$ at a distance of one from the vertical line that is part of any optimal solution. We also place $W$ points on each vertical line corresponding to picking a vertex at distance one from the horizontal line that is part of any solution and can now adjust the previous construction such that all points are covered by $2\ell\nu$ vertical lines, that is, we no longer need to imagine that points placed because of edges in the input graph (or for pairs of vertices in the same set $V_{i}$) have distinct distances. See fig. 3 for an illustration of the updated construction.
@@
Figure 3: The updated construction (for the same instance as in fig. 2). The violet lines represent lines that are part of any optimal solution and the blue dots with numbers next to them represent a collection of points at the same position. The value of $d_{s}$ is at least $500$. The colored dashed lines (together with the violet lines) represent an optimal solution corresponding to the independent set $\{a,e\}$. The horizontal line representing vertex $a$ is the closest line in the solution to $4$ points each on the horizontal lines representing $b$ and $c$. Since the distance to points on the line for $b$ is one, this adds a cost of four to the budget. Points on the line for $c$ have distance 22, so each points contributes 44. Summed up over all of these points, this gives an additional budget of $20$ which is balanced out by the blue vertex on the line representing $a$. For the edge representing $e$, there are 8 points at distance one (four on the lines for $d$ and $f$, respectively). Note that the blue point on the line for $e$ represents 88 points. Figure not drawn to scale due to the orders of magnitude difference in the distances.
@@
Finally, we want to ensure that lines in the solution are horizontal and/or vertical. Unfortunately, we were not able to ensure this property but the following construction ensures that all solution lines are almost vertical or horizontal. Consider the construction in fig. 4 where $d_{\ell}$ is much larger than $d_{s}$. We will show that any line that covers as many points as there are points in one row or column of the constructed point set, has to contain the points of exactly one row or one column. We will then pick a suitable number $p$, duplicate each point in the construction in fig. 3 $p$ times, and set the budget $B$ correspondingly such that any line that encodes a vertex in some set $V_{i}$ has a sum of squares of distances to points in the $i$th row or column of the construction in fig. 4 that is negligible while any line that is not almost vertical or horizontal has a point in any row or column of the construction such that the distance between this point and the line alone is larger than the budget $B$. We also place a lot of vertices in the eight “corners” of the construction to ensure that there are two horizontal and two vertical lines that are part of any solution. This concludes the overview of the reduction. We now present the construction in more detail.
@@
We set $k=2\ell+4$ and $B=n^{7}+(n-\ell)W+\ell\sum_{i=1}^{\nu}(W+\varphi(i))-\ell W+\ell p(n-\nu+1-q-\ell)$. Note that $B\leq n^{32}$. We then construct a set $F$ of $8n^{90}+k^{2}+2k$ points as follows. We place points in a $\frac{k}{2}\times(k+2)$ grid $G_{h}$ where the vertical distance between points is $d_{s}$ and the horizontal distance is $d_{\ell}$. We leave $(\ell+1)d_{s}$ extra space between columns $\nicefrac{{k+2}}{{2}}$ and $\nicefrac{{k+2}}{{2}}+1$ (this is where we will place the main part of the construction depicted in fig. 3). We place $n^{90}$ points at each of the four corners of the grid (on top of the already existing point). We then repeat the process with a $(k+2)\times\frac{k}{2}$ grid $G_{v}$ where the vertical distance between points is $d_{\ell}$ and the horizontal distance is $d_{s}$. We place the two grids such that the additional extra space forms a square in the middle of both grids. We call the set of all points placed so far $F$ and the vertical/horizontal lines that each go exactly through $2n^{90}+k+2$ points the fixed lines (we will show later that we can assume these four lines to be part of any optimal solution). An illustration of the points in $F$ is given in Figure 4.
@@
4 W[2]-hardness of covering points with subspaces
In this section, we show that Hyperplane Cover is $\operatorname{W}[2]$-hard when parameterized by $k$ even if $B=0$. To this end, we present a parameterized reduction from Dominating Set parameterized by solution size $k^{\prime}$. In Dominating Set, one is given a graph $G=(V,E)$ and an integer $k^{\prime}$ and the question is whether there exists a subset $S$ of vertices such that $S$ intersects with the closed neighborhood of every vertex. Dominating Set parameterized by $k^{\prime}$ is known to be W[2]-hard [cygan2015parameterized]. The basic idea of the following reduction is to capture the closed neighborhood of a vertex through an affine subspace.

### v2

@@ Abstract
Our main contributions highlight notable differences in the parameterized complexity of these problems. While Line Cover is fixed-parameter tractable with respect to the number $k$ of lines in the solution, we show that Line Clustering is $\operatorname{W}[1]$-hard parameterized by $k$, and that, under the Exponential Time Hypothesis, it admits no algorithm running in time $n^{o(k)}$. Moreover, although Hyperplane Cover has been known to be $\operatorname{NP}$-hard since the 1980s, following work of Megiddo and Tamir, even for $d=2$, we show that it remains NP-hard even when $k=2$.
We complement our hardness results by presenting an algorithm for Projective Clustering. We show that this problem is solvable in time $n^{\mathcal{O}(dk(r+1))}$. This not only yields an upper bound for Line Clustering that asymptotically matches our lower bound, but also significantly extends the benchmark exact algorithms for $k$-means clustering due to Inaba, Katoh, and Imai.
@@ 1 Introduction
The Line Cover problem asks for the smallest number $k$ of lines that cover a given set of $n$ points in the Euclidean plane $\mathbb{R}^{2}$. This problem is NP-hard and is well-studied in the context of parameterized complexity, particularly because it admits a simple kernelization, often highlighted in introductory lectures on kernelization [16, 25, 40].
@@
Line Clustering is an optimization variant of Line Cover, in which the goal is to find $k$ lines that best fit the points in the $\ell_{2}$ norm, rather than merely covering them. In other words, we search for $k$ lines such that the sum of the squares of distances from each point to the closest point on one of the lines is minimized.
@@
Figure 1 provides an example of Line Clustering with $k=3$. In Operations Research, this problem is known as Line Facility Location  [39]. The problem has been studied for convex and piecewise linear functions [7] as well as in the weighted setting [13].
Figure 1: An instance of Line Clustering with $k=3$ and a solution (the three colored lines). The points are colored based on the color of the nearest line in the selected solution. The dashed lines show the region where points have equal distance from two solution lines.
@@
Langerman and Morin [37] initiated the study of the parameterized complexity of Hyperplane Cover (as part of a more general algorithmic framework). They showed that the problem is FPT when parameterized by $k+d$, providing several algorithms running in $k^{\mathcal{O}(dk)}n^{\mathcal{O}(1)}$ time.
@@
Projective Clustering naturally finds many applications in fields such as unsupervised learning [42], data mining [3, 41], artificial intelligence [51], computational biology [42], database management [10], computer vision [43], image segmentation [52], motion segmentation [48], face clustering [29], image processing [30], systems theory [47], and others [23, 49]. It also generalizes a number of fundamental and well-studied problems such as $k$-Means Clustering [38], Principal Component Analysis (PCA[21], and Hyperplane Approximation [34].
@@
We consider 1 to be particularly interesting for several reasons. First, it underscores a clear contrast between Line Cover, which is $\operatorname{FPT}$ when parameterized by $k$, and Line Clustering. Second, Line Clustering is closely related to $k$-means clustering, where instead of fitting $k$ lines, one seeks $k$ points (i.e., zero-dimensional subspaces) that best fit the data in the $L_{2}$ norm. Notably, the parameterized complexity of $k$-means clustering in $\mathds{R}^{2}$ (with parameter $k$) remains a longstanding open question [15].
Our second lower bound concerns Hyperplane Cover. Since in $\mathds{R}^{2}$ this problem is equivalent to Line Cover, which is $\operatorname{NP}$-complete, we know that the problem is $\operatorname{Para-NP}$-complete when parameterized by $d$. On the other hand, by the work of Langerman and Morin [37], Hyperplane Cover is $\operatorname{FPT}$ when parameterized by $k$ and $d$. 2 refines this complexity picture by proving that Hyperplane Cover is $\operatorname{NP-Hard}$ even when $k=2$ (and $d$ is large compared to $k$).
@@
Hyperplane Cover is $\operatorname{NP}$-hard when $k=2$.
@@
Since $k$-Means Clustering is the (very) special case of Projective Clustering with $r=0$, 3 significantly extends the work of Inaba et al. [33], who gave an algorithm for $k$-Means Clustering running in $n^{\mathcal{O}(kd)}$ time. It is also worth noting that the tools we use to prove 3 are quite different from those employed by Inaba et al., who relied on weighted Voronoi diagrams. In contrast, the proof of 3 leverages a fundamental result from computational algebraic geometry, which computes a set of sample points that meets every realizable sign condition of a family of polynomials restricted to an algebraic set.
@@
Line Cover is NP-hard and APX-hard [4, 9, 39]. A number of results in parameterized complexity and kernelization in the literature is devoted to Line Cover and Hyperplane Cover. Langerman and Morin [37] proposed an $k^{dk+d}n^{\mathcal{O}(1)}$-time algorithm for Hyperplane Cover. Later, the running time was improved by Wang et al. [50] to $\tfrac{k^{(d-1)k}}{1.3^{k}}n^{\mathcal{O}(1)}$. Afshani et al. [1] gave an algorithm of running time $2^{n}\cdot n^{\mathcal{O}(1)}$. They also improved on the $\operatorname{FPT}$ algorithm for Hyperplane Cover in $\mathbb{R}^{3}$. Their improved algorithm runs in $\left(\tfrac{Ck^{2}}{\log k^{1/5}}\right)^{k}n^{\mathcal{O}(1)}$ time. They also presented a $\left(\tfrac{Ck}{\log k}\right)^{k}n^{\mathcal{O}(1)}$-time algorithm for Line Cover. Kratsch et al. show that the well-known $\mathcal{O}(k^{2})$ kernel for Line Cover is tight under standard complexity assumptions [35]. A similar result exists for Hyperplane Cover [8].
Due to its importance in theory as well as applications, the study of Projective Clustering enjoys a long and rich history. Various approaches have been used to solve Projective Clustering: kernel sets [2, 28], coresets [28, 46], total sensitivity [24, 36, 46], and dimension reduction via sampling [17, 44]. There is also a significant body of work on coresets constructions [19, 27, 31] and dimension reductions [11] for Projective Clustering in computational geometry.
@@ 2 Preliminaries
It is widely believed that $\operatorname{FPT}\neq\operatorname{W}[1]$, and hence, if a problem is hard for a class $\operatorname{W}[i]$ with $i\geq 1$, then it is considered to be fixed-parameter intractable. For a detailed introduction to parameterized complexity, we refer readers to the textbooks by Cygan et al. and by Downey and Fellows [16, 20]. We also provide conditional lower bounds by making use of the following complexity hypothesis formulated by Impagliazzo, Paturi, and Zane [32].
@@ 3 W[1]-hardness of projective clustering in the Euclidean plane
We give a reduction from Regular Multicolored Independent Set. Therein, one is given a graph $G=(V,E)$ and a partition of the $n$ vertices into $V_{1},V_{2},\ldots,V_{\ell}$ (often referred to as colors of vertices) such that $|V_{i}|=\nu=\nicefrac{{n}}{{\ell}}$ for each $i\in[\ell]$ and each vertex $v$ has exactly $q$ neighbors (none of which have the same color as $v$). The question is whether there exists an independent set containing exactly one vertex of each color. It is known that Regular Multicolored Independent Set is $\operatorname{W}[1]$-hard parameterized by $\ell$ [16, Proposition 13.5.] and cannot be solved in $n^{o(\ell)}$ time unless the $\operatorname{ETH}$ fails [12]. We will assume without loss of generality that $\nu$ is divisible by $4$, $\nu>\ell^{3}$, $\ell>10$, and $V_{i}=\{w_{1}^{i},w_{2}^{i},\ldots,w_{\nu}^{i}\}$.
@@
In the remaining part of the proof, we provide the gap-reduction for a multiset of input points. Before we present the reduction in detail, we first give a high-level overview. Our aim is to build a grid-like structure where any optimal solution contains $\ell$ vertical and $\ell$ horizontal lines. Each horizontal line and each vertical line correspond to picking a vertex in a solution. The aim is then to construct a set of points such that (i) for each horizontal line in the solution there is a vertical line in the solution that encodes the same vertex (to ensure that a solution selects precisely $\ell$ vertices) and (ii) for each pair of horizontal and vertical lines in a solution, the corresponding vertices are not adjacent in the input graph. We do this by placing points on the intersection of two such lines whenever the two respective vertices share an edge or are two distinct vertices of the same color. Otherwise, we place a point on the horizontal line close to the intersection. Imagine for now that these points have distinct distances from a “solution line” such that no vertical line contains more than one such point. See Figure 2 for an example.
@@
With this intuition in mind, there are two main obstacles to overcome. First, we want to ensure that the solution lines are actually (almost) horizontal and vertical and second, covering a maximum number of points is different from minimizing the sum of squares of distances from each point to the closest line. We start with the second obstacle. Assume that there are two lines (one vertical and one horizontal) that is part of any optimal solution that is far enough away from the previously described set of points so that they are never the closest line to any of these points. Let $W$ be some sufficiently large (but also not too large) number. We place all vertical lines corresponding to vertices within the same set $V_{i}$ closer together than lines corresponding to vertices in different sets (as already indicated in Figure 2). We do the same for all horizontal lines. By making the distance $d_{s}$ between lines corresponding to vertices in different sets large enough, we can ensure that for any set $V_{i}$ any optimal solution contains one horizontal line and one vertical line that corresponds to a vertex in $V_{i}$ (assuming that all solution lines are horizontal or vertical). Next, we compute for each horizontal line the sum of squares of distances of all points on horizontal lines corresponding to vertices within the same set $V_{i}$ except for those points that lie on the vertical line corresponding to the same vertex. For a line $h$, let that number be $\phi(h)$. Then, we place $W-\phi(h)$ points on line $h$ at a distance of one from the vertical line that is part of any optimal solution. We also place $W$ points on each vertical line corresponding to picking a vertex at distance one from the horizontal line that is part of any solution and can now adjust the previous construction such that all points are covered by $2\ell\nu$ vertical lines, that is, we no longer need to imagine that points placed because of edges in the input graph (or for pairs of vertices in the same set $V_{i}$) have distinct distances. See Figure 3 for an illustration of the updated construction.
@@
Figure 3: The updated construction (for the same instance as in Figure 2). The violet lines represent lines that are part of any optimal solution and the blue dots with numbers next to them represent a collection of points at the same position. The value of $d_{s}$ is at least $500$. The colored dashed lines (together with the violet lines) represent an optimal solution corresponding to the independent set $\{a,e\}$. The horizontal line representing vertex $a$ is the closest line in the solution to $4$ points each on the horizontal lines representing $b$ and $c$. Since the distance to points on the line for $b$ is one, this adds a cost of four to the budget. Points on the line for $c$ have distance 22, so each points contributes 44. Summed up over all of these points, this gives an additional budget of $20$ which is balanced out by the blue vertex on the line representing $a$. For the edge representing $e$, there are 8 points at distance one (four on the lines for $d$ and $f$, respectively). Note that the blue point on the line for $e$ represents 88 points. Figure not drawn to scale due to the orders of magnitude difference in the distances.
@@
Finally, we want to ensure that lines in the solution are horizontal and/or vertical. Unfortunately, we were not able to ensure this property but the following construction ensures that all solution lines are almost vertical or horizontal. Consider the construction in Figure 4 where $d_{\ell}$ is much larger than $d_{s}$. We will show that any line that covers as many points as there are points in one row or column of the constructed point set, has to contain the points of exactly one row or one column. We will then pick a suitable number $p$, duplicate each point in the construction in Figure 3 $p$ times, and set the budget $B$ correspondingly such that any line that encodes a vertex in some set $V_{i}$ has a sum of squares of distances to points in the $i$th row or column of the construction in Figure 4 that is negligible while any line that is not almost vertical or horizontal has a point in any row or column of the construction such that the distance between this point and the line alone is larger than the budget $B$. We also place a lot of vertices in the eight “corners” of the construction to ensure that there are two horizontal and two vertical lines that are part of any solution. This concludes the overview of the reduction. We now present the construction in more detail.
@@
We set $k=2\ell+4$ and $B=n^{7}+(n-\ell)W+\ell\sum_{i=1}^{\nu}(W+\varphi(i))-\ell W+\ell p(n-\nu+1-q-\ell)$. Note that $B\leq n^{32}$. We then construct a set $F$ of $8n^{90}+k^{2}+2k$ points as follows. We place points in a $\frac{k}{2}\times(k+2)$ grid $G_{h}$ where the vertical distance between points is $d_{s}$ and the horizontal distance is $d_{\ell}$. We leave $(\ell+1)d_{s}$ extra space between columns $\nicefrac{{k+2}}{{2}}$ and $\nicefrac{{k+2}}{{2}}+1$ (this is where we will place the main part of the construction depicted in Figure 3). We place $n^{90}$ points at each of the four corners of the grid (on top of the already existing point). We then repeat the process with a $(k+2)\times\frac{k}{2}$ grid $G_{v}$ where the vertical distance between points is $d_{\ell}$ and the horizontal distance is $d_{s}$. We place the two grids such that the additional extra space forms a square in the middle of both grids. We call the set of all points placed so far $F$ and the vertical/horizontal lines that each go exactly through $2n^{90}+k+2$ points the fixed lines (we will show later that we can assume these four lines to be part of any optimal solution). An illustration of the points in $F$ is given in Figure 4.
@@
4 $\operatorname{NP}$-hardness of covering points with two subspaces
In this section, we show that Hyperplane Cover is $\operatorname{NP}$-hard even when $k=2$. To this end, we present a reduction from Equal-Cardinality Number Partition (ECNP, in short). In ECNP, we are given a set $W=\{w_{1},\dots,w_{2n}\}$ of $2n$ positive integers that sum to $2t$, and the question is whether $W$ can be partitioned into two disjoint subsets $R_{1}$ and $R_{2}$ such that $|R_{1}|=|R_{2}|=n$, and $\sum_{w\in R_{1}}w=\sum_{w\in R_{2}}w=t$. ECNP is known to be $\operatorname{NP}$-hard [26].
Let Equal-Cardinality Number Partition with distinct numbers (ECNP-d, in short) denote the restriction of the problem where the numbers in the input are distinct. It is a folklore knowledge that ECNP-d is $\operatorname{NP}$-hard (the result can be shown by disturbing the integers in the input of ECNP). The basic idea of the following reduction is to capture the $n$ distinct numbers that add up to $t$ via a hyperplane.